A MOBILE ROBUST FTIR FOR PASSIVE FIELD DETERMINATION OF GASEOUS SPECIES

R. DAUM, A. BEIL, Bruker Saxonia Analytik, Permoserstraße 15 Leipzig, D-04318 Germany; \underline{T. J. JOHNSON}, B. M. ABRAHAM, Bruker Optics, 19 Fortune Drive, Billerica, MA 01821.

Passive remote sensing using an FTIR spectrometer allows the detection and identification of pollutant/toxic gases in the atmosphere. In this paper we report a measurement method and data reduction technique that to not require a previously measured reference spectrum. Recent experimental results obtained using a new high-sensitivity mobile FTIR are presented. After a radiometric calibration of the spectrometer using reference blackbodies, the spectral radiance of the background is determined. Using the inverse function of Planck’s radiation law, the brightness temperature ϵ(ν) is computed. The temperature spectrum provides a spectrally constant reference for many of the materials used as backdrops in field measurements since their emittance is high and nearly constant in the 800 to 1200 cm^{−1} range. Experimental results are presented to illustrate the enhancement of the signal-to-noise ratio that can be achieved by using a backdrop that provides maximal temperature difference to ambient environment.